40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 4 16:43:50 2014, manasa, Update, General, IFO alignment update 
    Reply  Tue Feb 4 17:40:19 2014, manasa, Update, General, Ready for pump down?? 
       Reply  Tue Feb 4 23:01:24 2014, Jenne, Update, General, Ready for pump down?? 
          Reply  Wed Feb 5 15:26:57 2014, manasa, Update, General, Ready for pump down?? 
       Reply  Mon Feb 10 16:02:12 2014, Jenne, Update, General, PRM oplev clipped in vacuum Screenshot-Untitled_Window-1.png
          Reply  Mon Feb 10 16:20:28 2014, Jenne, Update, General, PRM QPD recentered 
             Reply  Mon Feb 10 18:59:25 2014, Jenne, Update, ASC, PRM ASC better, but not great yet 
                Reply  Mon Feb 10 19:56:10 2014, rana, Update, ASC, PRM ASC better, but not great yet 
Message ID: 9596     Entry time: Tue Feb 4 16:43:50 2014     Reply to this: 9597
Author: manasa 
Type: Update 
Category: General 
Subject: IFO alignment update 

[EricQ, Manasa]

We are close to the end of the vent except for a couple of issues.

* POP is not visible on the IR card. But we see POP flashes unclipped on the camera and also spikes in POP DC. So  we are assuming that the POP path hasn't gone far off. If anybody has suggestions for a better method to check this, we could give it a try.

* PRM suspension has not been behaving well. PRM is being kicked around every 5-10 seconds when the PRC is aligned (as seen on REFL camera). We are not sure where this is coming from. The first time we saw this happening was when we were trying to lock PRC at low power even before we took the heavy doors off. So we are pretty sure this is not caused by the foil cover on the OSEMs. We tried turning ON/OFF the oplev servo, turning ON/OFF the damping loops and also checked the connections in the feedthrough and satellite box for the PRM. The OSEM sensor values for the suspension also seem to match the ones on the wiki.

 

Closeup checklist

 

  • Center beam on all AS optics
  • GET CAMERA IMAGES OF EVERYTHING

    • We must get images right before closing, right after closing, etc.
  • Make sure REFL is clear
    • dither PRM, see motion on AP tables
  • Make sure AS is clear
    • dither BS/ITM, see motion on AP tables
  • Using IPANG/POS pick-off mirrors, center beams on:
    • IPPOS
    • IPANG (IPANG aligned to be low in pitch. The beam comes out of vacuum unclipped and hits the first steering mirror outside vacuum at its lower edge)
  • Check green alignment in the arms and make sure the transmitted green reaches the PSL table.
  • Check all OpLevs centered, in and out of vacuum

  • Close PSL shutter & green shutters at the ends

  • Check jam nuts
ELOG V3.1.3-